OpenStudy (anonymous):

what are three equivalent ratios for 20:40

OpenStudy (australopithecus):

this is just a fraction, \[20:40 = \frac{20}{40}\]

OpenStudy (australopithecus):

Do you know how to simplify fractions?

OpenStudy (australopithecus):

what number can you divide the numerator and denominator by to give you whole numbers, for example \[\frac{4}{6} = \frac{\frac{4}{2}}{\frac{6}{2}} = \frac{2}{3}\]

OpenStudy (australopithecus):

notice 2/3 cannot be divided by any one number to give a whole number on the numerator or denominator, thus it is in its simplest form

OpenStudy (australopithecus):

Just say you wanted to represent a fraction as something larger, you can also multiply the numerator and denominator by any one number and still obtain the same fraction, this is useful for addition and subtraction of fractions. For example, \[\frac{2}{3} = \frac{2}{2}*\frac{2}{3} = \frac{2*2}{2*3} = \frac{4}{6}\] Notice that, \[\frac{2}{2} = 1\]

OpenStudy (australopithecus):

so you aren't changing the value of the fraction just the way it is being represented

OpenStudy (australopithecus):

Hope this helps good luck and if you have any questions feel free to ask

OpenStudy (anonymous):

10:20 30:60 1000:2000

OpenStudy (australopithecus):

@Madhdi2020 this isn't an answer site

OpenStudy (australopithecus):

Explain the concepts at least and show how you got those answers, otherwise you are potentially helping someone cheat on their homework. This stuff is important and the person asking the question needs to learn it.